# BhashaVerse : Translation Ecosystem for Indian Subcontinent Languages

### v1

@@
12024
[…]
BhashaVerse : Translation Ecosystem for Indian Subcontinent Languages
@@ 1 Introduction
Addressing these challenges requires a systematic approach. Developing translation systems for Indian languages is crucial to facilitate communication and preserve linguistic diversity. With the translation system, robust evaluation frameworks are needed to assess their quality, with or without reference translations. Identifying and categorizing translation errors is essential for system improvement, enabling targeted refinements based on predefined error groups. Automatic post-editing systems can then help correct translation errors efficiently, either independently or in collaboration with human translators. Incorporating these advancements across general and domain-specific applications will greatly benefit the language research community and end-users.
@@
Finally, this work addresses several translation-related applications that are central to enhancing multilingual communication.
@@
    Discourse Translation: Translation at the discourse level addresses coherence, context retention, and inter-sentence dependencies. Our work evaluates how well current systems handle these aspects and proposes improvements for translating connected text effectively. Discourse translation possible for 3636 language pairs with keeping 1536 tokens as a context with encoder-decoder model while 4096 context for decoder only model.
@@
    Domain-Specific Translation: Different domains, such as healthcare, legal, and education, require specialized vocabulary and context-aware translations. We design systems capable of handling domain-specific language nuances to improve usability in practical scenarios. We include domains
@@ 3 Developed Corpora : Machine Translation
3.1 Automatic Alignment and Human Validation
@@
3.2 Human Post-edited Education and Medical Domain Parallel Corpora
@@
3.3 Quality Synthetic Parallel Corpora
@@
3.3.1 Pivoted Parallel Corpora
@@
3.3.2 Iterative Backward Translation
We employed back-translation, a widely used data augmentation technique in machine translation, to enhance the diversity and quality of training data. Monolingual corpora for the target languages were sourced from WMT News Crawl [^21] https://data.statmt.org/news-crawl/ and Wikipedia [^22] https://huggingface.co/datasets/wikimedia/wikipedia. Using SSMT Translator [^23] http://ssmt.iiit.ac.in/translatev3 and IndicTrans2242424https://github.com/AI4Bharat/IndicTrans2, we translated this monolingual data bidirectionally: from Indian languages to English and from English to Indian languages. The translated data introduced variations in sentence structure, vocabulary, and phrasing, enriching the training set’s diversity. For each monolingual corpus, both translation outputs were evaluated using COMET-QE scores, and the version with the highest score was retained. This filtered data was then used to train a multilingual machine translation model involving the selected language pairs. The newly trained model was subsequently used to re-translate the same monolingual corpora, iteratively improving the translations. This process was repeated for five iterations, with the best-scored translation pairs from all iterations being selected for the final training corpus.
@@ 3.3.2 Iterative Backward Translation
For the Wikipedia dataset [^25] https://huggingface.co/datasets/wikimedia/wikipedia and crawled news articles from the internet, we adopted paragraph-level backward translation to preserve coherence and contextual consistency across entire paragraphs. Unlike sentence-level back-translation, this approach ensures that the translated content retains discourse-level attributes, such as topic continuity, pronoun resolution, and lexical cohesion. By focusing on paragraph-level fidelity, this method enhances the quality of machine translation systems, particularly for tasks requiring deeper contextual understanding, such as document translation or conversational AI. The same filtering mechanism was applied here as well. Each sentence within a paragraph was evaluated for quality using COMET-QE scores, and the best-scored translations were selected. These high-quality sentence translations were then reassembled into paragraphs to maintain the original discourse structure. This strategy ensured the development of a robust multilingual machine translation model capable of handling complex and contextually rich translations.
3.4 Parallel Corpora Cleaning
@@
The quality of source-target and target-source sentence pairs was evaluated using COMET-22 scores. Sentence pairs with scores significantly below the average threshold (average minus 10) were excluded, ensuring that only high-quality data was retained for training.
@@
4 Automatic Post-Editing (APE) Corpora
4.1 Human APE Corpora
As discussed in Section 3.2, post-editing involves human translators refining machine-generated translations to produce accurate and natural-sounding final translations. This process was conducted using the PostEditMe [^26] https://posteditme.in/ workbench, a platform designed to facilitate post-editing tasks. The workbench allowed translators to edit machine-generated translations while maintaining a record of both the original machine-generated version and the human-corrected version. Additionally, the workbench captured all intermediate stages of post-editing, recording multiple iterations for each sentence pair, along with the final edited version. These intermediate stages and the final post-edited text form a comprehensive resource for studying the post-editing process and its impact on translation quality. The resulting dataset, named the Bhashik APE corpus, includes the initial machine-generated (also intermediate stages text) and human-post-edited translations for source texts in English and 11 Indian languages. This corpus supports the development of automatic post-editing systems and enhances domain-specific machine translation by addressing the linguistic nuances of technical and educational contexts more effectively.
4.2 Synthetically Developed Corpora
Automatic post-editing (APE) corpora typically consist of three key components: the source text, an imperfect or erroneous translation, and the error-free translation. To synthetically generate such APE corpora, we utilized both existing parallel corpora and our newly developed parallel corpora described in this work. The target sentences or passages are subjected to various perturbation techniques as detailed in Appendix Section 7, simulating errors commonly found in machine translations. For our opted language pairs, we processed each sentence pair to create multiple denoised versions of the target text. These perturbations introduced varying degrees of errors, ranging from 2% to 15% of the total words, including grammar mistakes, mistranslations, and word order issues. These artificially generated error patterns simulate typical MT errors, making the perturbed version function as the erroneous translation, while the original target text serves as the error-free translation. Additionally, we applied noisy back-translation techniques to create diverse post-editing pairs, ensuring a wide range of error types for robust APE training data. This approach provides a rich and diverse dataset for training automatic post-editing models. We call the resulting dataset Bhashik APE (Synth).
5 Corpora for Machine Translation Evaluation
5.1 Human Annotated Corpora
To develop these corpora, we utilized benchmark datasets and multiple translation engines, including Google Translate, SSMTv3, and Indictransv2, to generate translations for input texts. Experts evaluated both the source text and the translated text, providing ratings. The reference translations in the benchmark corpora were used for reference-enabled machine translation evaluation, while the annotations can also be applied in reference-free evaluation scenarios. The dataset spans 18 language pairs and 36 translation directions, resulting in over 72,000 rated translations. This effort involved over 70 experienced language experts, with each source and target translation pair receiving three independent ratings.
5.2 Synthetically Developed Corpora
  1. 1.
  2. 2.
    • Source text (S)
    • Original translation text (T) from the parallel corpora
    • Perturbed translation text (T̃) with Error percentage (Ep)
  3. 3.
    Calculating Synthetic Scores: Using the existing COMET Quality Estimation (QE) model, we first evaluated the source text (S) and the original translation (T) to obtain a COMET score (Cs) and 100 minus the Translation Error Rate (TER) score (Tp) using original translation (T) and Perturbed translation text (T̃). We then averaged the error percentage (Ep) and the TER-derived score (Tp) to calculate a degradation factor. This factor was subtracted from the COMET score (Cs) to derive a new synthetic score for the source text (S) and the perturbed translation (T̃).
By applying this process, we generated an extensive synthetic dataset to supplement the limited human-annotated data, enabling robust evaluation across a broader range of translation quality scenarios. The resulting dataset is named the Bhashik PseudoMTDA corpus for English and 11 Indian languages.
6 Corpora for Machine Translation Error Span Identification and Error Category Marking
Error identification and categorization help analyze weaknesses in MT systems, enabling targeted improvements. This section outlines our development of corpora to support these tasks.
6.1 Human MT Error Span Identification Corpora
As discussed in Section 3.2, post-editing involves human translators refining machine-generated translations to produce accurate, natural-sounding final translations. This process was carried out using the PostEditMe [^27] https://posteditme.in/ workbench, a platform specifically designed to support post-editing tasks. The workbench captures every edit made by translation experts and records this information at the word level over the machine translation output. The edited spans in the translated text, where changes were made, are marked as MT error spans for the corresponding translation. These post-editing efforts provide a rich resource for identifying error spans and evaluating translation quality. The resulting dataset, named the Bhashik PseudoMTError corpus, includes the source text, the initial machine-generated translation, and the annotated error spans for English and 11 Indian languages.
6.2 Synthetically Developed MT Error Span and Category Marking Corpora
For the selected language pairs, we processed each sentence pair to generate multiple denoised versions of the target text. Perturbations introduced errors ranging from 2% to 15% of the total words, including grammar mistakes, mistranslations, and word order issues. These artificially generated error patterns mimic typical MT errors, making the perturbed versions serve as erroneous translations. Error spans were annotated based on the introduced errors, along with the corresponding categories used to generate the perturbations.
This approach provides a rich and diverse dataset for training automatic systems to mark MT error spans and categories. We refer to the resulting dataset as the Bhashik PseudoMTError Corpus.
@@ 6.3 Developed Parallel Corpora
6.3.1 Bhashik Education and Medical: Human Post-edited Corpora
Table 1: The post-edited 2M domain-parallel corpora spans English and 10 Indian languages, covering 15 educational domains.
Law eng_Latn-tel_Telu 80165 Marketing eng_Latn-tel_Telu 4971
Law eng_Latn-mar_Deva 81581 Marketing eng_Latn-mar_Deva 4858
Law eng_Latn-tam_Taml 90 Marketing eng_Latn-kan_Knda 5099
Mathematics eng_Latn-tam_Taml 286 Textile eng_Latn-kan_Knda 7630
Mathematics eng_Latn-mal_Mlym 163 History eng_Latn-tel_Telu 4902
health eng_Latn-tam_Taml 12728 History eng_Latn-mar_Deva 4919
health eng_Latn-guj_Gujr 299 History eng_Latn-kan_Knda 4941
health eng_Latn-tel_Telu 34855 Economics eng_Latn-tel_Telu 5087
health eng_Latn-hin_Deva 11802 Economics eng_Latn-mar_Deva 5093
health eng_Latn-mar_Deva 1904 Economics eng_Latn-kan_Knda 5061
health eng_Latn-kan_Knda 135 Tourism eng_Latn-tel_Telu 12798
health eng_Latn-mal_Mlym 202 Tourism eng_Latn-mar_Deva 12795
health eng_Latn-ben_Beng 370 Tourism eng_Latn-kan_Knda 12813
Teaching methods eng_Latn-mar_Deva 32327
6.3.2 Bhashik Generic : Quality Synthetic Corpora
Table 2: The generated 1B parallel corpora cover 325 language pairs, encompassing 36 languages from the Indian subcontinent and utilizing 15 unique scripts.
@@ 6.3.2 Bhashik Generic : Quality Synthetic Corpora
7 Text Perturbation
We employ various text perturbation techniques to generate task-specific corpora for translation-related applications. By leveraging existing parallel and monolingual corpora, we introduce controlled modifications using carefully designed methods. These denoising strategies, which include random additions, deletions, and substitutions, effectively simulate errors in spelling, grammar, word order, punctuation, and more for all 36 languages mentioned in this work. The crafted perturbation methods are outlined below:

### v2

@@
Abstract
We present a multilingual, multi-task translation model for 36 Indian languages, including Assamese, Hindi, Tamil, and Urdu. The model utilizes 10 billion parallel corpora for 36 x 36 language directions and millions of data pairs for tasks like translation evaluation, post-editing, grammar correction, and error identification. Corpus creation involves leveraging existing resources, synthetic data, and domain-specific datasets. The model’s multi-task approach integrates translation, grammar correction, error identification, quality estimation, and post-editing, ensuring robust cross-lingual communication. Evaluations cover discourse-level and domain-specific translation, using reference-based and reference-free metrics. This work enhances translation quality, supports linguistic diversity, and advances low-resource language processing.
@@ 1 Introduction
Addressing these challenges requires a systematic approach. Developing translation systems for Indian languages is crucial to facilitate communication and preserve linguistic diversity. With the translation system, robust evaluation frameworks are needed to assess their quality, with or without reference translations. Identifying and categorizing translation errors is essential for system improvement, enabling targeted refinements based on predefined error groups. Automatic post-editing systems can then help correct translation errors efficiently, either independently or in collaboration with human translators. Incorporating these advancements across general and domain-specific applications will greatly benefit the language research community and end users.
@@
Finally, this work addresses the following translation-related tasks that are central to enhancing multilingual communication in the Indian subcontinent.
@@
    Discourse Translation: Translation at the discourse level addresses coherence, context retention, and inter-sentence dependencies. Our work evaluates how well current systems handle these aspects and proposes improvements for translating connected text effectively. Discourse translation possible for 36 x 36 language pairs with keeping 1536 tokens as a context with encoder-decoder model while 4096 context for decoder only model.
@@
    Domain-Specific Translation: Different domains, such as healthcare, legal, and education, require specialized vocabulary and context-aware translations. We design systems capable of handling domain-specific language nuances to improve usability in practical scenarios.
@@ 3 Developed Corpora : Machine Translation
Automatic Alignment and Human Validation
@@
Human Post-edited Education and Medical Domain Parallel Corpora
@@
Quality Synthetic Parallel Corpora
@@
Pivoted Parallel Corpora
@@
Iterative Backward Translation
We employed back-translation, a widely used data augmentation technique in machine translation, to enhance the diversity and quality of training data. Monolingual corpora for the target languages were sourced from WMT News Crawl [^21] https://data.statmt.org/news-crawl/ and Wikipedia [^22] https://huggingface.co/datasets/wikimedia/wikipedia. Using SSMT Translator [^23] http://ssmt.iiit.ac.in/translatev3 and IndicTrans2242424https://github.com/AI4Bharat/IndicTrans2, we translated this monolingual data bidirectionally: from Indian languages to English and from English to Indian languages. The translated data introduced variations in sentence structure, vocabulary, and phrasing, enriching the training set’s diversity. For each monolingual corpus, both translation outputs were evaluated using COMET-QE scores, and the version with the highest score is retained. This filtered data is then used to train a multilingual machine translation model involving the selected language pairs. The newly trained model is subsequently used to re-translate the same monolingual corpora, iteratively improving the translations. This process is repeated for five iterations, with the best-scored translation pairs from all iterations being selected for the final training corpus.
@@
For the Wikipedia dataset [^25] https://huggingface.co/datasets/wikimedia/wikipedia and crawled news articles from the internet and from the available monolingual corpora [^26] https://data.statmt.org/news-crawl/, we adopted paragraph-level backward translation to preserve coherence and contextual consistency across entire paragraphs. Unlike sentence-level back-translation, this approach ensures that the translated content retains discourse-level attributes, such as topic continuity, pronoun resolution, and lexical cohesion. By focusing on paragraph-level fidelity, this method enhances the quality of machine translation systems, particularly for tasks requiring deeper contextual understanding, such as document translation or conversational AI. The same filtering mechanism is applied here as well. Each sentence within a paragraph is evaluated for quality using COMET-QE scores, and the best-scored translations were selected. These high-quality sentence translations were then reassembled into paragraphs to maintain the original discourse structure. This strategy ensured the development of a robust multilingual machine translation model capable of handling complex and contextually rich translations.
Parallel Corpora Cleaning
@@
The quality of source-target and target-source sentence pairs is evaluated using COMET-22 scores. Sentence pairs with scores significantly below the average threshold (average minus 10) are excluded, ensuring that only high-quality data is retained for training.
@@ 3.1 Developed Parallel Corpora
Table 1: The generated over 1B parallel corpora cover 325 language pairs, encompassing 36 languages from the Indian subcontinent and utilizing 15 unique scripts.
@@
Table 2: Postedited Health Domain Corpora for English and 8 Indian Languages
Language Pair Passages/Sentences Language Pair Passages/Sentences
ben_Beng_Latn-tam_Taml 142 eng_Latn-tel_Telu 10308
ben_Beng_Latn-tel_Telu 291 hin_Deva-ben_Beng_Latn 6384
eng_Latn-ben_Beng_Latn 7915 hin_Deva-guj_Gujr 145
eng_Latn-guj_Gujr 158 hin_Deva-kan_Knda 1138
eng_Latn-hin_Deva 11042 hin_Deva-mal_Mlym 1436
eng_Latn-kan_Knda 1209 hin_Deva-mar_Deva 879
eng_Latn-mal_Mlym 1754 hin_Deva-tam_Taml 5517
eng_Latn-mar_Deva 906 hin_Deva-tel_Telu 6412
eng_Latn-tam_Taml 8726 tam_Taml-tel_Telu 1019
Table 3: The human post-edited 2M domain-parallel corpora spans English and 10 Indian languages, covering 15 educational domains.
Law eng_Latn-tel_Telu 80165 Marketing eng_Latn-tel_Telu 4971
Law eng_Latn-mar_Deva 81581 Marketing eng_Latn-mar_Deva 4858
Law eng_Latn-tam_Taml 90 Marketing eng_Latn-kan_Knda 5099
Mathematics eng_Latn-tam_Taml 286 Textile eng_Latn-kan_Knda 7630
Mathematics eng_Latn-mal_Mlym 163 History eng_Latn-tel_Telu 4902
health eng_Latn-tam_Taml 12728 History eng_Latn-mar_Deva 4919
health eng_Latn-guj_Gujr 299 History eng_Latn-kan_Knda 4941
health eng_Latn-tel_Telu 34855 Economics eng_Latn-tel_Telu 5087
health eng_Latn-hin_Deva 11802 Economics eng_Latn-mar_Deva 5093
health eng_Latn-mar_Deva 1904 Economics eng_Latn-kan_Knda 5061
health eng_Latn-kan_Knda 135 Tourism eng_Latn-tel_Telu 12798
health eng_Latn-mal_Mlym 202 Tourism eng_Latn-mar_Deva 12795
health eng_Latn-ben_Beng 370 Tourism eng_Latn-kan_Knda 12813
Teaching methods eng_Latn-mar_Deva 32327
In this section, we introduce the corpora we developed for machine translation between English and 35 Indian languages, covering both general and domain-specific contexts, as detailed in Tables 1, 3 and 2. These parallel corpora are publicly available for research purposes on Hugging Face  [^27] https://huggingface.co/ltrciiith under the names: bhashik-parallel-corpora-generic, bhashik-parallel-corpora-education, and bhashik-parallel-corpora-health.
4 Automatic Post-Editing (APE) Corpora
Table 4: Automatic post-editing corpora for English and Indian languages: general domain (Bhashik APE Synth - 10M) and education and health domains (Bhashik APE Human - 2M).
Bhashik APE (Synth) : General Bhashik APE (Synth) : General Bhashik APE (Human) : Education and Health
Source Language Target Language #Sentences Source Language Target Language #Sentences Source Language Target Language #Sentences
eng_Latn san_Deva 167912 asm_Beng eng_Latn 307940 eng_Latn tam_Taml 13958
eng_Latn mni_Mtei 75170 ben_Beng eng_Latn 498159 eng_Latn mar_Deva 733974
eng_Latn ory_Orya 188394 brx_Deva eng_Latn 221311 eng_Latn kan_Knda 664516
eng_Latn doi_Deva 95643 doi_Deva eng_Latn 197047 eng_Latn hin_Deva 407093
eng_Latn mai_Deva 124677 gom_Deva eng_Latn 201763 eng_Latn urd_Arab 4307
eng_Latn tam_Taml 113849 guj_Gujr eng_Latn 327686 eng_Latn tel_Telu 751077
eng_Latn mar_Deva 297651 hin_Deva eng_Latn 564166 eng_Latn mal_Mlym 999
eng_Latn kan_Knda 135478 kan_Knda eng_Latn 432813
eng_Latn npi_Deva 231413 kas_Arab eng_Latn 196620
eng_Latn hin_Deva 261516 kas_Deva eng_Latn 43189
eng_Latn urd_Arab 225262 mai_Deva eng_Latn 156935
eng_Latn asm_Beng 258234 mal_Mlym eng_Latn 524830
eng_Latn guj_Gujr 146769 mar_Deva eng_Latn 274688
eng_Latn brx_Deva 126492 mni_Beng eng_Latn 42669
eng_Latn ben_Beng 264087 mni_Mtei eng_Latn 203056
eng_Latn gom_Deva 88346 npi_Deva eng_Latn 561283
eng_Latn tel_Telu 192549 ory_Orya eng_Latn 385591
eng_Latn sat_Olck 43932 pan_Guru eng_Latn 293203
eng_Latn kas_Arab 45590 san_Deva eng_Latn 217712
eng_Latn mal_Mlym 723 sat_Olck eng_Latn 139586
tel_Telu eng_Latn 465576 snd_Arab eng_Latn 24459
urd_Arab eng_Latn 457598 snd_Deva eng_Latn 93589
tam_Taml eng_Latn 312192
Human APE Corpora
As discussed in Section 3, post-editing involves human translators refining machine-generated translations to produce accurate and natural-sounding final translations. This task was conducted using the PostEditMe [^28] https://posteditme.in/ workbench, a platform designed to facilitate post-editing tasks. The workbench allowed translators to edit machine-generated translations while maintaining a record of both the original machine-generated version and the human-corrected version. Additionally, the workbench captured all intermediate stages of post-editing, recording multiple iterations for each sentence pair, along with the final edited version. These intermediate stages and the final post-edited text form a comprehensive resource for studying the post-editing process and its impact on translation quality. The resulting dataset, named the Bhashik APE corpus (presented in Table 4), includes the initial machine-generated (also intermediate stages text) and human-post-edited translations for source texts in English and 11 Indian languages. This corpus supports the development of automatic post-editing systems and enhances domain-specific machine translation by addressing the linguistic nuances of technical and educational contexts more effectively.
Synthetically Developed Corpora
Automatic post-editing (APE) corpora typically consist of three key components: the source text, an imperfect or erroneous translation, and the error-free translation. To synthetically generate such APE corpora, we utilized both existing parallel corpora and our newly developed parallel corpora described in this work. The target sentences or passages are subjected to various perturbation techniques as detailed in Appendix Section 7, simulating errors commonly found in machine translations. For our opted language pairs, we processed each sentence pair to create multiple denoised versions of the target text. These perturbations introduced varying degrees of errors, ranging from 2% to 15% of the total words, including grammar mistakes, mistranslations, and word order issues. These artificially generated error patterns simulate typical MT errors, making the perturbed version function as the erroneous translation, while the original target text serves as the error-free translation. Additionally, we applied noisy back-translation techniques to create diverse post-editing pairs, ensuring a wide range of error types for robust APE training data. This approach provides a rich and diverse dataset for training automatic post-editing models. We call the resulting dataset Bhashik APE (Synth) and details are presented in Table 4.
5 Corpora for Machine Translation Evaluation
Table 5: 1M Human direct assessment corpora (without reference) for English and Indian languages across general domain for 36 language pairs.
Source Language Target Language #Sentences Source Language Target Language #Sentences
eng_Latn ben_Beng 2960 hin_Deva ben_Beng 3001
eng_Latn guj_Gujr 2961 hin_Deva guj_Gujr 2961
eng_Latn hin_Deva 2961 hin_Deva kan_Knda 3001
eng_Latn kan_Knda 3001 hin_Deva mar_Deva 3201
eng_Latn mar_Deva 3001 hin_Deva ory_Orya 2399
eng_Latn ory_Orya 2361 hin_Deva pan_Guru 3000
eng_Latn Panjabi 2961 hin_Deva tel_Telu 3041
eng_Latn tam_Taml 2961 hin_Deva urd_Arab 2053
eng_Latn tel_Telu 2961 ben_Beng hin_Deva 3001
eng_Latn urd_Arab 1994 guj_Gujr hin_Deva 2961
ben_Beng eng_Latn 2960 kan_Knda hin_Deva 3001
guj_Gujr eng_Latn 2961 mar_Deva hin_Deva 3201
hin_Deva eng_Latn 2961 ory_Orya hin_Deva 2080
kan_Knda eng_Latn 3001 pan_Guru hin_Deva 3000
mar_Deva eng_Latn 3001 tel_Telu hin_Deva 3041
ory_Orya eng_Latn 2321 urd_Arab hin_Deva 2001
pan_Guru eng_Latn 2961 tel_Telu eng_Latn 2961
tam_Taml eng_Latn 2961 urd_Arab eng_Latn 1961
Human Annotated Corpora
To develop these corpora, we utilized benchmark datasets and multiple translation engines, including Google Translate, SSMTv3, and Indictransv2, to generate translations for input texts. Experts evaluated both the source text and the translated text, providing ratings. The reference translations in the benchmark corpora were used for reference-enabled machine translation evaluation, while the annotations can also be applied in reference-free evaluation scenarios. The dataset spans 18 language pairs and 36 translation directions, resulting in over 101K rated translations. This effort involved over 70 experienced language experts, with each source and target translation pair receiving three independent ratings. The human-annotated corpora details are presented in Table 5.
Table 6: 10M pseudo direct assessment corpora (with reference) for English and Indian languages across general, educational, and health domains.
Source Language Target Language #Sentences Source Language Target Language #Sentences
asm_Beng eng_Latn 275688 eng_Latn hin_Deva 668620
ben_Beng eng_Latn 298946 eng_Latn mar_Deva 1031650
brx_Deva eng_Latn 139431 eng_Latn tel_Telu 943632
doi_Deva eng_Latn 113633 eng_Latn urd_Arab 229569
gom_Deva eng_Latn 112863 eng_Latn kan_Knda 800024
guj_Gujr eng_Latn 154981 eng_Latn tam_Taml 127947
hin_Deva eng_Latn 356329 eng_Latn ben_Beng 264087
kan_Knda eng_Latn 354733 eng_Latn asm_Beng 258234
kas_Arab eng_Latn 138187 eng_Latn ory_Orya 188394
kas_Deva eng_Latn 42465 eng_Latn mai_Deva 124677
mai_Deva eng_Latn 148104 eng_Latn brx_Deva 126492
mal_Mlym eng_Latn 258292 eng_Latn doi_Deva 95643
mar_Deva eng_Latn 159852 eng_Latn npi_Deva 231413
mni_Beng eng_Latn 42324 eng_Latn guj_Gujr 146769
mni_Mtei eng_Latn 121557 eng_Latn mni_Mtei 75170
npi_Deva eng_Latn 285402 eng_Latn san_Deva 167912
ory_Orya eng_Latn 206850 eng_Latn gom_Deva 88346
pan_Guru eng_Latn 37968 eng_Latn kas_Arab 45590
san_Deva eng_Latn 171657 eng_Latn sat_Olck 43932
sat_Olck eng_Latn 139451 eng_Latn mal_Mlym 1722
snd_Deva eng_Latn 62937 tel_Telu eng_Latn 331194
tam_Taml eng_Latn 129059 urd_Arab eng_Latn 255991
Table 7: 70M pseudo direct assessment corpora (without reference) for English and Indian languages across general domain.
Source Language Target Language #Sentences Source Language Target Language #Sentences
asm_Beng eng_Latn 615562 eng_Latn ben_Beng 3050599
ben_Beng eng_Latn 101559 eng_Latn hin_Deva 4949515
bho_Deva eng_Latn 2991 eng_Latn guj_Gujr 316079
brx_Deva eng_Latn 25297 eng_Latn mar_Deva 214706
doi_Deva eng_Latn 115211 eng_Latn pan_Guru 13973
gom_Deva eng_Latn 23858 eng_Latn kan_Knda 56693
guj_Gujr eng_Latn 307059 eng_Latn tam_Taml 10747282
hin_Deva eng_Latn 11386159 eng_Latn tel_Telu 307867
kan_Knda eng_Latn 63748 eng_Latn mag_Deva $92$
kas_Arab eng_Latn 149232 eng_Latn mal_Mlym 8963878
kas_Deva eng_Latn 202814 eng_Latn npi_Deva 51496
lus_Latn eng_Latn 3021 eng_Latn kas_Deva 33387
mai_Deva eng_Latn 92319 eng_Latn brx_Deva 50430
mal_Mlym eng_Latn 8883930 eng_Latn san_Deva 277518
mar_Deva eng_Latn 75266 eng_Latn kas_Arab 149251
mni_Beng eng_Latn 8775 eng_Latn gom_Deva 27831
mni_Mtei eng_Latn 25508 eng_Latn sat_Olck 25126
npi_Deva eng_Latn 1635019 eng_Latn mni_Mtei 25471
ory_Orya eng_Latn 2959267 eng_Latn asm_Beng 109159
pan_Guru eng_Latn 27787 eng_Latn bho_Deva 2955
san_Deva eng_Latn 277532 eng_Latn ory_Orya 2936801
sat_Olck eng_Latn 25140 eng_Latn lus_Latn 3006
sin_Sinh eng_Latn 3045 eng_Latn mai_Deva 30065
snd_Arab eng_Latn 5688 eng_Latn sin_Sinh 3012
snd_Deva eng_Latn 13114 eng_Latn mni_Beng 8855
tam_Taml eng_Latn 10614897 eng_Latn snd_Arab 5575
tel_Telu eng_Latn 351303 eng_Latn urd_Arab 46907
urd_Arab eng_Latn 47896 eng_Latn doi_Deva 24177
eng_Latn hoc_Wara 201 eng_Latn snd_Deva 13027
Synthetically Developed Corpora
  1. 1.
  2. 2.
    • Source text (S)
    • Original translation text (T) from the parallel corpora
    • Perturbed translation text (T̃) with Error percentage (Ep)
  3. 3.
    Calculating Synthetic Scores: Using the existing COMET Quality Estimation (QE) model, we first evaluated the source text (S) and the original translation (T) to obtain a COMET score (Cs) and 100 minus the Translation Error Rate (TER) score (Tp) using original translation (T) and Perturbed translation text (T̃). We then averaged the error percentage (Ep) and the TER-derived score (Tp) to calculate a degradation factor. This factor is subtracted from the COMET score (Cs) to derive a new synthetic score for the source text (S) and the perturbed translation (T̃).
By applying this process, we generated an extensive synthetic dataset to supplement the limited human-annotated data, enabling robust evaluation across a broader range of translation quality scenarios. The resulting dataset is named the Bhashik PseudoMTDA corpus for English and 11 Indian languages. The developed corpora for direct assessment, both with and without reference, are shown in Table 6 and Table 7, respectively.
6 Corpora for MT Error Span Identification
Table 8: 20M pseudo error-annotated corpora with 54M error spans for English and Indian languages across general, educational, and health domains.
Source Language Target Language #Sentences #Errors #Errors per Sent Source Language Target Language #Sentences #Errors #Errors per Sent
asm_Beng eng_Latn 551376 1278030 2.32 eng_Latn hin_Deva 1337218 4849894 3.63
ben_Beng eng_Latn 597892 1385958 2.32 eng_Latn ben_Beng 528174 1205310 2.28
brx_Deva eng_Latn 278862 633800 2.27 eng_Latn mar_Deva 2063250 8725642 4.23
doi_Deva eng_Latn 227266 508742 2.24 eng_Latn npi_Deva 727670 926702 1.27
gom_Deva eng_Latn 225726 493676 2.19 eng_Latn asm_Beng 516468 1237170 2.4
guj_Gujr eng_Latn 309962 705820 2.28 eng_Latn san_Deva 337358 1875910 5.56
hin_Deva eng_Latn 712636 1225358 1.72 eng_Latn mni_Mtei 150340 479422 3.19
kan_Knda eng_Latn 707896 1022196 1.44 eng_Latn urd_Arab 459138 816608 1.78
kas_Arab eng_Latn 276374 663646 2.4 eng_Latn mai_Deva 249354 597684 2.4
kas_Deva eng_Latn 84930 251870 2.97 eng_Latn brx_Deva 252984 684588 2.71
mai_Deva eng_Latn 296208 665144 2.25 eng_Latn tel_Telu 1887252 6017654 3.19
mal_Mlym eng_Latn 516584 1189960 2.3 eng_Latn kas_Arab 141542 244316 1.73
mni_Beng eng_Latn 84648 251446 2.97 eng_Latn ory_Orya 379626 869170 2.29
mni_Mtei eng_Latn 243114 539906 2.22 eng_Latn tam_Taml 270474 375414 1.39
npi_Deva eng_Latn 570804 1317958 2.31 eng_Latn doi_Deva 191286 481650 2.52
ory_Orya eng_Latn 413792 942118 2.28 eng_Latn guj_Gujr 293538 620304 2.11
pan_Guru eng_Latn 75936 158936 2.09 eng_Latn kan_Knda 1599988 5099336 3.19
san_Deva eng_Latn 343314 774460 2.26 eng_Latn gom_Deva 353456 296182 0.84
sat_Olck eng_Latn 278902 620176 2.22 eng_Latn sat_Olck 87864 201198 2.29
snd_Deva eng_Latn 125874 277914 2.21 eng_Latn mal_Mlym 311942 3244 0.01
tam_Taml eng_Latn 258118 584962 2.27 tel_Telu eng_Latn 661054 954970 1.44
urd_Arab eng_Latn 511982 1183818 2.31
Error identification and categorization help analyze weaknesses in MT systems, enabling targeted improvements. This section outlines our development of corpora to support these tasks. As discussed in Section 3, post-editing involves human translators refining machine-generated translations to produce accurate, natural-sounding final translations. This task was carried out using the PostEditMe [^29] https://posteditme.in/ workbench, a platform specifically designed to support post-editing tasks. The workbench captures every edit made by translation experts and records this information at the word level over the machine translation output. The edited spans in the translated text, where changes were made, are marked as MT error spans for the corresponding translation. These post-editing efforts provide a rich resource for identifying error spans and evaluating translation quality. The resulting dataset, named the Bhashik PseudoMTError corpus, includes the source text, the initial machine-generated translation, and the annotated error spans for English and multiple Indian languages.
Synthetically Developed MT Error Span Marked Corpora
For the selected language pairs, we processed each sentence pair to generate multiple denoised versions of the target text. Perturbations introduced errors ranging from 2% to 15% of the total words, including grammar mistakes, mistranslations, and word order issues. These artificially generated error patterns mimic typical MT errors, making the perturbed versions serve as erroneous translations. Error spans were annotated based on the introduced errors to generate the perturbations.
This approach provides a rich and diverse dataset for training automatic systems to mark MT error spans and categories. We refer to the resulting dataset as the Bhashik PseudoMTError Corpus and details are presented in Table 8.
7 Corpora for Grammar Correction
We use text perturbation techniques to generate grammar correction corpora by introducing controlled errors in monolingual data. These include spelling, grammar, word order, and punctuation errors for all 36 languages. The methods are outlined below: